Why momentum squared is zero for photons?
1
Expert's answer
2019-06-17T14:05:33-0400

For relativistic particles, dispersion relation reads


"p_\mu p^\mu = \frac{E^2}{c^2}-\vec{p}^2=m^2c^2"

where "p_\mu" is momentum 4 vector, "E" is energy of particle, "\vec{p}" is momentum , m is the mass and c is speed of light.

For photons, mass equals to zero - because they travel with maximum possible speed c.

Therefore

"p_\mu p^\mu=0"